# Are all models wrong?Fundamental limits in distribution-free empirical model falsification

### v1

@@ 5 Discussion
In this paper, we have discussed a central question in learning theory: when can we show that a model class is “wrong”, in the sense that even the best instance of a model in it cannot achieve zero test error? We do so by considering the more general task of lower bounding the model class risk empirically and in a distribution-free manner. Focusing on the case of highly flexible model classes capable of often perfectly interpolating the data, we identify a regime of “mild” interpolation (where, although $\mathcal{F}$ interpolates the data $\mathcal{D}_{n}$, it is not able to interpolate $\gg n^{2}$ many data points), a setting in which we may still empirically be able to perform nontrivial inference on $R_{P}(\mathcal{F})$. In contrast, in the high-complexity setting (where $\mathcal{F}$ can interpolate $\gg n^{2}$ many data points), we may say that we are in a regime of “hyper-interpolation”, where nontrivial inference is no longer possible. Our examples show how these regimes are sharply characterized.
@@ 5.1 Related literature
As discussed in Section 1, the model class risk is closely related to the question of excess risk and the approximation–estimation tradeoff (see, for example, Bartlett et al., (2002); Bottou and Bousquet, (2007)). Specifically, we have seen that the problem of finding an upper bound on excess risk is essentially equivalent in difficulty to the problem of a lower bound for $R_{P}(\mathcal{F})$—and consequently, the hardness results established in our work can be interpreted as evidence that a distribution-free upper bound on excess risk is also challenging in a regime where the model class $\mathcal{F}$ is highly complex.
@@
In recent years, an exciting topic in machine learning research has been the benign overfitting phenomenon, sometimes also called double descent, where an increase in a model class’s capacity beyond the interpolation threshold can help to reduce generalization error of the eventually fitted function (Belkin et al., 2019a, ). In other words, fitting a model $\hat{f}$ with $\hat{R}(\hat{f},\mathcal{D}_{n})=0$ (meaning, implicitly, that we are searching within in a class $\mathcal{F}$ with $N(\mathcal{F},P)\geq n$) can often lead to high accuracy, in contrast to what earlier results, based on generalization bounds from the learning theory literature, might suggest. There are a number of analyses for the generalization error upper bound for different overfitting estimators; see for example Belkin et al., (2018); Belkin et al., 2019b ; Bartlett et al., (2020).
@@
thus providing a distribution-free upper bound in the sense of Definition 1. Such high-probability upper confidence bounds are the starting point for the distribution-free risk control methodology of Bates et al., (2021); see also Angelopoulos et al., (2022, 2023) for related approaches.
@@ 5.2 Extensions and open questions
Second, recall that in this work we have identified an “in-between” regime: as summarized in Figure 1, if the interpolation capacity of $\mathcal{F}$ lies between $\mathcal{O}(n)$ and $\mathcal{O}(n^{2})$, then it may be possible or impossible to provide distribution-free inference on $R_{P}(\mathcal{F})$. Our two examples (in Section 4) show that these boundaries cannot be improved—no universal guarantee of nontrivial inference can be achieved outside of the low-complexity regime, and no universal hardness result can be shown outside of the high-complexity regime. However, an important open question remains: might there be an alternative notion of model class complexity (i.e., different from the interpolation capacity), with which it would be possible to identify a universal phase transition directly from the low-complexity to high-complexity regime?
@@ Acknowledgments
R.F.B. was partially supported by the National Science Foundation via grant DMS-2023109, and by the Office of Naval Research via grant N00014-24-1-2544.
@@ A.1 A tighter lower bound for the low-complexity regime
Fix $\alpha\in(0,1)$, $n\geq 1$, and model class $\mathcal{F}$. Assume that the loss function $\ell$ is bounded, taking values in $[0,B]$. Define $\Delta_{n}\in(0,1]$ as the unique solution of [^5] Note that, since $\Delta\mapsto-\Delta-\log(1-\Delta)$ is strictly increasing, with $\lim_{\Delta\searrow 0}\{-\Delta-\log(1-\Delta)\}=0$ and $\lim_{\Delta\nearrow 1}\{-\Delta-\log(1-\Delta)\}=+\infty$, this is well-defined for any value $\hat{R}(\mathcal{F},\mathcal{D}_{n})\geq 0$. In particular if $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$, then this equation is solved by $\Delta_{n}=1$.
@@
$\hat{L}_{\alpha}^{\mathrm{ERM},B}(\mathcal{F},\mathcal{D}_{n})\geq\biggl{(}1-\sqrt{\frac{2B\log(1/\alpha)}{n\hat{R}(\mathcal{F},\mathcal{D}_{n})}}\biggr{)}\hat{R}(\mathcal{F},\mathcal{D}_{n})=\hat{R}(\mathcal{F},\mathcal{D}_{n})-\sqrt{\hat{R}(\mathcal{F},\mathcal{D}_{n})\cdot\frac{2B\log(1/\alpha)}{n}}.$
@@
In the case that $R_{P}(\mathcal{F})=0$, we have $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$ almost surely, and so the result holds trivially. From this point on, then, we will assume $R_{P}(\mathcal{F})>0$ (note that we still may have $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$).
First, for any fixed $f\in\mathcal{F}$, note that $\hat{R}(f,\mathcal{D}_{n})$ is an average of $n$ i.i.d. terms, which each lie in $[0,B]$ and have mean $R_{P}(f)$. By a multiplicative Chernoff bound (McDiarmid,, 1998, Theorem 2.3(b)), we therefore have
$\mathbb{P}\biggl{\{}\hat{R}(f,\mathcal{D}_{n})\leq(1+\delta)R_{P}(f)\biggr{\}}\geq 1-\exp\left(-h(\delta)\cdot\frac{nR_{P}(f)}{B}\right)$
for $\delta>0$, where $h(\delta)=(1+\delta)\log(1+\delta)-\delta$. Next, for any $\varepsilon>0$, we can find some $f_{\varepsilon}\in\mathcal{F}$ such that $R_{P}(\mathcal{F})\leq R_{P}(f_{\varepsilon})\leq R_{P}(\mathcal{F})+\varepsilon$. Applying the above bound, then,
$\mathbb{P}\biggl{\{}\hat{R}(f_{\varepsilon},\mathcal{D}_{n})\leq(1+\delta)\cdot\big{(}R_{P}(\mathcal{F})+\varepsilon\big{)}\biggr{\}}\geq 1-\exp\left(-h(\delta)\cdot\frac{nR_{P}(\mathcal{F})}{B}\right).$
Since $\hat{R}(f_{\varepsilon},\mathcal{D}_{n})\geq\inf_{f\in\mathcal{F}}\hat{R}(f,\mathcal{D}_{n})=\hat{R}(\mathcal{F},\mathcal{D}_{n})$, therefore,
$\mathbb{P}\biggl{\{}\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq(1+\delta)\cdot\big{(}R_{P}(\mathcal{F})+\varepsilon\big{)}\biggr{\}}\geq 1-\exp\left(-h(\delta)\cdot\frac{nR_{P}(\mathcal{F})}{B}\right).$
And, since $\varepsilon>0$ can be chosen to be arbitrarily small, it therefore holds that
$\mathbb{P}\biggl{\{}\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq(1+\delta)R_{P}(\mathcal{F})\biggr{\}}\geq 1-\exp\left(-h(\delta)\cdot\frac{nR_{P}(\mathcal{F})}{B}\right).$
Next, let $\delta_{n}>0$ be the unique solution to
$h(\delta_{n})=\frac{B\log(1/\alpha)}{nR_{P}(\mathcal{F})},$
so that we have
$\mathbb{P}\biggl{\{}\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq(1+\delta_{n})R_{P}(\mathcal{F})\biggr{\}}\geq 1-\alpha.$
To complete the proof, we therefore only need to verify that
$\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq(1+\delta_{n})R_{P}(\mathcal{F})\ \Longrightarrow\ (1-\Delta_{n})\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F}).$
First, define
$\Delta_{n}^{\prime}=\frac{\delta_{n}}{1+\delta_{n}}\in(0,1).$
Then $1+\delta_{n}=1/(1-\Delta^{\prime}_{n})$, and so
$\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq(1+\delta_{n})R_{P}(\mathcal{F})\ \Longrightarrow\ (1-\Delta^{\prime}_{n})\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F}).$
Our last step is to check that, on the event $(1-\Delta^{\prime}_{n})\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F})$, we have $\Delta^{\prime}_{n}\leq\Delta_{n}$. We have
$\frac{B\log(1/\alpha)}{nR_{P}(\mathcal{F})}=h(\delta_{n})=(1+\delta_{n})\log(1+\delta_{n})-\delta_{n}=\frac{1}{1-\Delta^{\prime}_{n}}\log\left(\frac{1}{1-\Delta^{\prime}_{n}}\right)-\frac{\Delta^{\prime}_{n}}{1-\Delta^{\prime}_{n}},$
so rearranging terms,
$-\Delta^{\prime}_{n}-\log(1-\Delta^{\prime}_{n})=(1-\Delta^{\prime}_{n})\cdot\frac{B\log(1/\alpha)}{nR_{P}(\mathcal{F})}\leq\frac{B\log(1/\alpha)}{n\hat{R}(\mathcal{F},\mathcal{D}_{n})}=-\Delta_{n}-\log(1-\Delta_{n}),$
where the inequality holds on the event that $(1-\Delta^{\prime}_{n})\hat{R}(\mathcal{F},\mathcal{D}_{n})\leq R_{P}(\mathcal{F})$, and the last step holds by definition of $\Delta_{n}$. Finally, since $\Delta\mapsto-\Delta-\log(1-\Delta)$ is a strictly increasing function, this implies $\Delta^{\prime}_{n}\leq\Delta_{n}$, which completes the proof.
@@ A.2 Proof of Theorem 4 (the hardness result)
The proof of this result follows the “sample–resample” strategy used for establishing certain hardness results in the distribution-free inference literature (see Angelopoulos et al., (2024, Lemma 4.14)), which relies on the fact that, when sampling $n$ times from a population of size $N$, the total variation distance between sampling with replacement and sampling without replacement is bounded by $\frac{n^{2}}{2N}$.
@@
by noting that $R_{Q}(\mathcal{F})=\hat{R}(\mathcal{F},(z_{1},\dots,z_{N}))$ by definition of $Q$ as an empirical distribution. Next, consider indices $J_{1},\dots,J_{n}\in[N]$ sampled uniformly without replacement. By the total variation distance bound mentioned above, we therefore have

### v2

@@ 5 Discussion
In this paper, we have discussed a central question in learning theory: when can we show that a model class is “wrong”, in the sense that even the best instance of a model in the class cannot achieve zero test error? We do so by considering the more general task of lower bounding the model class risk empirically and in a distribution-free manner. Focusing on the case of highly flexible model classes capable of often perfectly interpolating the data, we identify a regime of “mild” interpolation (where, although $\mathcal{F}$ interpolates the data $\mathcal{D}_{n}$, it is not able to interpolate $\gg n^{2}$ many data points), a setting in which we may still empirically be able to perform nontrivial inference on $R_{P}(\mathcal{F})$. In contrast, in the high-complexity setting (where $\mathcal{F}$ can interpolate $\gg n^{2}$ many data points), we may say that we are in a regime of “hyper-interpolation”, where nontrivial inference is no longer possible. Our examples show how these regimes are sharply characterized.
@@ 5.1 Related literature
As discussed in Section 1, the model class risk is closely related to the question of excess risk $R_{P}(\hat{f})-\inf_{f\in\mathcal{F}}R_{P}(f)$ and the approximation–estimation tradeoff (see, for example, Bartlett et al., (2002); Bottou and Bousquet, (2007)). Specifically, the problem of finding an upper bound on excess risk is essentially equivalent in difficulty to the problem of a lower bound for $R_{P}(\mathcal{F})$—and consequently, the hardness results established in our work can be interpreted as evidence that a distribution-free upper bound on excess risk is also challenging in a regime where the model class $\mathcal{F}$ is highly complex.
@@
In recent years, an exciting topic in machine learning research has been the benign overfitting phenomenon, sometimes also called double descent, where an increase in a model class’s capacity beyond the interpolation threshold can help to reduce generalization error of the eventually fitted function (Belkin et al., 2019a, ). In other words, fitting a model $\hat{f}$ with $\hat{R}(\hat{f},\mathcal{D}_{n})=0$ (meaning, implicitly, that we are searching within in a class $\mathcal{F}$ with $N(\mathcal{F},P)\geq n$) can often lead to high accuracy, in contrast to what earlier results, based on generalization bounds from the learning theory literature, might suggest. There are a number of analyses for the generalization error upper bound for different overfitting estimators; see for example Belkin et al., (2018); Belkin et al., 2019b ; Bartlett et al., (2020). A few recent works have also provided lower bounds on the excess risk for estimators with large training errors to illustrate that interpolation or memorization is necessary for optimal generalization, for example, see Cheng et al., (2022) and references therein.
@@
thus providing a distribution-free upper bound in the sense of Definition 1. Such high-probability upper confidence bounds are the starting point for the distribution-free risk control methodology of Bates et al., (2021); see also Angelopoulos et al., (2025, 2023) for related approaches.
An important existing result on distribution-free inference on a model class risk is given by Devroye et al., (1996, Theorem 8.5), which illustrates that in a distribution-free classification setting, the task of estimating the Bayes error leads to a constant minimax lower bound.
@@ 5.2 Extensions and open questions
Second, recall that in this work we have identified an “in-between” regime: as summarized in Figure 1, if the interpolation capacity of $\mathcal{F}$ lies between $n$ and $\mathcal{O}(n^{2})$, then it may be possible or impossible to provide distribution-free inference on $R_{P}(\mathcal{F})$. Our two examples (in Section 4) show that these boundaries cannot be improved—no universal guarantee of nontrivial inference can be achieved outside of the low-complexity regime, and no universal hardness result can be shown outside of the high-complexity regime. However, an important open question remains: might there be an alternative notion of model class complexity (i.e., different from the interpolation capacity), with which it would be possible to identify a universal phase transition directly from the low-complexity to high-complexity regime?
@@ Acknowledgments
R.F.B. was partially supported by the National Science Foundation via grant DMS-2023109, and by the Office of Naval Research via grant N00014-24-1-2544. We thank John Duchi for bringing helpful references to our attention.
@@ A.1 A tighter lower bound for the low-complexity regime
Fix $\alpha\in(0,1)$, $n\geq 1$, and model class $\mathcal{F}$. Assume that the loss function $\ell$ is bounded, taking values in $[0,B]$ for some $B>0$. Define $\Delta_{n}\in(0,1]$ as the unique solution of [^7] Note that, since $\Delta\mapsto-\Delta-\log(1-\Delta)$ is strictly increasing on $[0,1)$, with $\lim_{\Delta\searrow 0}\{-\Delta-\log(1-\Delta)\}=0$ and $\lim_{\Delta\nearrow 1}\{-\Delta-\log(1-\Delta)\}=+\infty$, this is well-defined for any value $\hat{R}(\mathcal{F},\mathcal{D}_{n})\geq 0$. In particular if $\hat{R}(\mathcal{F},\mathcal{D}_{n})=0$, then this equation is solved by $\Delta_{n}=1$.
@@
$\hat{L}_{\alpha}^{\mathrm{ERM},B}(\mathcal{F},\mathcal{D}_{n})\geq\biggl{(}1-\sqrt{\frac{2B\log(1/\alpha)}{n\hat{R}(\mathcal{F},\mathcal{D}_{n})}}\biggr{)}\hat{R}(\mathcal{F},\mathcal{D}_{n})=\hat{R}(\mathcal{F},\mathcal{D}_{n})-\sqrt{\hat{R}(\mathcal{F},\mathcal{D}_{n})\cdot\frac{2B\log(1/\alpha)}{n}}.$ (A.2)
@@
In fact, this type of result can be applied to the setting of an unbounded loss, as well, via a truncation step. For any $f\in\mathcal{F}$, define its truncated empirical risk as $\hat{R}(f,\mathcal{D}_{n};B)=\frac{1}{n}\sum_{i=1}^{n}\min\{\ell(f,Z_{i}),B\}$, and define $\hat{R}(\mathcal{F},\mathcal{D}_{n};B)=\inf_{f\in\mathcal{F}}\hat{R}(f,\mathcal{D}_{n};B)$.
Theorem A.2.
Fix $\alpha\in(0,1)$, $n\geq 1$, and model class $\mathcal{F}$. Fix any $B>0$, and define $\Delta_{n,B}\in(0,1]$ as the unique solution of
$-\Delta_{n,B}-\log(1-\Delta_{n,B})=\frac{B\log(1/\alpha)}{n\hat{R}(\mathcal{F},\mathcal{D}_{n};B)}.$
Then
$\hat{L}_{\alpha}^{\mathrm{ERM-trunc},B}(\mathcal{F},\mathcal{D}_{n}):=(1-\Delta_{n,B})\hat{R}(\mathcal{F},\mathcal{D}_{n};B)$
is a valid distribution-free lower bound on $R_{P}(\mathcal{F})$.
To help interpret this lower bound, we can observe that as in (A.2) above, it holds that
$\hat{L}_{\alpha}^{\mathrm{ERM-trunc},B}(\mathcal{F},\mathcal{D}_{n})\geq\hat{R}(\mathcal{F},\mathcal{D}_{n};B)-\sqrt{\hat{R}(\mathcal{F},\mathcal{D}_{n};B)\cdot\frac{2B\log(1/\alpha)}{n}}\\ \geq\hat{R}(\mathcal{F},\mathcal{D}_{n};B)-\sqrt{\frac{2B^{2}\log(1/\alpha)}{n}},$ (A.3)
where the last step holds simply because $\hat{R}(\mathcal{F},\mathcal{D}_{n};B)\leq B$ by construction.
The proofs of these results rely on the following lemma.
Lemma A.3.
Let $Z_{1},\ldots,Z_{n}$ be independent random variables taking values in $[0,1]$. Denote $S:=\sum_{i=1}^{n}Z_{i}$ and $\mu:=\mathbb{E}[S]$. Fix $\alpha\in(0,1)$ and let $\Delta\in(0,1]$ be the unique solution to
$-\Delta-\log(1-\Delta)=\log(1/\alpha)/S.$
Then
$\mathbb{P}\bigl{(}(1-\Delta)S\leq\mu\bigr{)}\geq 1-\alpha.$
Proof of Lemma A.3.
First we consider the degenerate case where $\mu=0$, i.e., $Z_{i}=0$ almost surely for each $i$. Then, almost surely, $S=0$ and $\Delta=1$, and the result holds trivially. From this point on, then, we assume $\mu>0$ (note that we might still have $S=0$ with positive probability).
By a multiplicative Chernoff bound (McDiarmid,, 1998, Theorem 2.3(b)), we have
$\mathbb{P}\Bigl{\{}S\leq(1+\delta)\mu\Bigr{\}}\geq 1-\exp\bigl{(}-h(\delta)\mu\bigr{)}$
for $\delta>0$, where $h(\delta)=(1+\delta)\log(1+\delta)-\delta$. Since $h$ is continuous and strictly increasing on $[0,\infty)$, with $h(0)=0$ and $h(\delta)\rightarrow\infty$ as $\delta\rightarrow\infty$, there exists a unique $\delta_{\mu}>0$ such that $h(\delta_{\mu})=\log(1/\alpha)/\mu$, so that we have
$\mathbb{P}\Bigl{\{}S\leq(1+\delta_{\mu})\mu\Bigr{\}}\geq 1-\alpha.$
To complete the proof, we therefore only need to verify that
$S\leq(1+\delta_{\mu})\mu\ \Longrightarrow\ (1-\Delta)S\leq\mu.$
First, define
$\Delta^{\prime}=\frac{\delta_{\mu}}{1+\delta_{\mu}}\in(0,1).$
Then $1+\delta_{\mu}=1/(1-\Delta^{\prime})$, and so
$S\leq(1+\delta_{\mu})\mu\ \Longrightarrow\ (1-\Delta^{\prime})S\leq\mu.$
Our last step is to check that, on the event $(1-\Delta^{\prime})S\leq\mu$, we have $\Delta^{\prime}\leq\Delta$. We have
$\frac{\log(1/\alpha)}{\mu}=h(\delta_{\mu})=(1+\delta_{\mu})\log(1+\delta_{\mu})-\delta_{\mu}=\frac{1}{1-\Delta^{\prime}}\log\left(\frac{1}{1-\Delta^{\prime}}\right)-\frac{\Delta^{\prime}}{1-\Delta^{\prime}},$
so rearranging terms,
$-\Delta^{\prime}-\log(1-\Delta^{\prime})=(1-\Delta^{\prime})\cdot\frac{\log(1/\alpha)}{\mu}\leq\frac{\log(1/\alpha)}{S}=-\Delta-\log(1-\Delta),$
where the inequality holds on the event that $(1-\Delta^{\prime})S\leq\mu$, and the last step holds by definition of $\Delta$. Finally, since $u\mapsto-u-\log(1-u)$ is a strictly increasing function on $[0,1)$, this implies $\Delta^{\prime}\leq\Delta$, which completes the proof. ∎
We are now in a position to state the proofs of Theorems A.1 and A.2.
@@
This result is simply a special case of Theorem A.2 (since, when the loss $\ell$ takes values in $[0,B]$, we have $\hat{R}(\mathcal{F},\mathcal{D}_{n};B)=\hat{R}(\mathcal{F},\mathcal{D}_{n})$). ∎
Proof of Theorem A.2.
For any $\varepsilon>0$, we can find some $f_{\varepsilon}\in\mathcal{F}$ such that $R_{P}(\mathcal{F})\leq R_{P}(f_{\varepsilon})\leq R_{P}(\mathcal{F})+\varepsilon$. Noting that $\frac{\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)}{B}$ is an average of $n$ i.i.d. terms, which each lie in $[0,1]$, we can apply Lemma A.3, by taking $\Delta_{\varepsilon}$ to be the unique solution to
$-\Delta_{\varepsilon}-\log(1-\Delta_{\varepsilon})=\frac{B\log(1/\alpha)}{n\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)}$
to see that
$1-\alpha\leq\mathbb{P}\biggl{\{}(1-\Delta_{\varepsilon})\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\biggr{\}}\\ \leq\mathbb{P}\biggl{\{}(1-\Delta_{\varepsilon})\hat{R}(\mathcal{F},\mathcal{D}_{n};B)\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\biggr{\}},$
where the last step holds since $\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\geq\hat{R}(\mathcal{F},\mathcal{D}_{n};B)$. And, again using the fact that $\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)\geq\hat{R}(\mathcal{F},\mathcal{D}_{n};B)$, we have
$-\Delta_{\varepsilon}-\log(1-\Delta_{\varepsilon})=\frac{B\log(1/\alpha)}{n\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)}\leq\frac{B\log(1/\alpha)}{n\hat{R}(\mathcal{F},\mathcal{D}_{n};B)}=-\Delta_{n,B}-\log(1-\Delta_{n,B})$
by definition of $\Delta_{n,B}$. Since $u\mapsto-u-\log(1-u)$ is strictly increasing on $[0,1)$, it follows that we must have $\Delta_{\varepsilon}\leq\Delta_{n,B}$. We conclude
$1-\alpha\leq\mathbb{P}\biggl{\{}(1-\Delta_{n,B})\hat{R}(\mathcal{F},\mathcal{D}_{n};B)\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\biggr{\}}.$
Next, we also have
$\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n};B)]\leq\mathbb{E}[\hat{R}(f_{\varepsilon},\mathcal{D}_{n})]=R_{P}(f_{\varepsilon})\leq R_{P}(\mathcal{F})+\varepsilon,$
where the first step holds since $\hat{R}(f,\mathcal{D}_{n};B)\leq\hat{R}(f,\mathcal{D}_{n})$ for any $f\in\mathcal{F}$ by definition of the truncated empirical risk, and the last step holds by definition of $f_{\varepsilon}$. Therefore,
$\mathbb{P}\biggl{\{}(1-\Delta_{n,B})\hat{R}(\mathcal{F},\mathcal{D}_{n};B)\leq R_{P}(\mathcal{F})+\varepsilon\biggr{\}}\geq 1-\alpha,$
and since $\varepsilon>0$ can be chosen to be arbitrarily small, the result follows. ∎
@@ A.2 Proof of Theorem 4 (the hardness result)
by noting that $R_{Q}(\mathcal{F})=\hat{R}(\mathcal{F},(z_{1},\dots,z_{N}))$ by definition of $Q$ as an empirical distribution. Next, consider indices $J_{1},\dots,J_{n}\in[N]$ sampled uniformly without replacement. By a total variation distance bound on the difference between sampling with and without replacement (e.g., Stam, (1978); see also Angelopoulos et al., (2024, Lemma 4.15)), we therefore have